Help me with this question please.
At present Mary is four younger than her brother Tom.If Tom is n years old,what will be the sum of their age in three years time?If the sum of their age is 15 then,how old will Tom be?

Answer by Alan3354(69443)

At present Mary is four younger than her brother Tom. If Tom is n years old,what will be the sum of their age in three years time?
--
The difference remains 4 years.
Tom is n, Mary is n-4.
In 3 years, Tom will be n+3, Mary will be n-1.
(n+3) + (n-1) = 2n+2
================
If the sum of their age is 15 then,how old will Tom be?
How old will he be when?
--------
Now?
T - M = 4
T + M = 15
------------- Add
2T = 19
Tom is 9.5
Mary is 5.5